# Recorded Versus Synthetic Ground Motions: A Comparative Analysis of Structural Seismic Responses

### v1

@@ 3.1 Metric 1: Probability distributions
Figures 5 and 6 present the empirical PDFs and complementary cumulative distribution functions (CCDFs) of EDPs—peak IDRs and PFAs—for the medium-period building and the long-period tower. Both structural models exhibit close agreement between responses to synthetic and recorded ground motions in the central ranges of the PDFs, indicating comparable global response characteristics under typical seismic scenarios. This alignment suggests that calibrating the synthetic and recorded ground motions to a target spectrum ensures consistency in global response characteristics. However, discrepancies become apparent in the tails of the distributions, particularly for the IDRs of the long-period tower. These differences can be attributed to the non-bijective mapping between spectral acceleration and ground motions, where both selection and generation methods introduce variability in time-domain characteristics despite comparable response spectra.
@@ 3.2 Metric 2: Statistical moments
Figure 8 illustrates differences in normalized moments ($DF_{M_{k}}$, $i=k,...,4$)—mean, variance, skewness, and kurtosis—between responses to synthetic and recorded ground motions for both structural archetypes. Differences in mean and variance remain within 3%–20%, indicating close agreement in central tendencies and overall variability. These results demonstrate that well-calibrated synthetic motions can effectively capture general response characteristics.
@@ 3.3 Metric 3: Correlations between EDPs
Figure 9 presents correlation matrices for EDPs under synthetic and recorded motions, along with the corresponding differences quantified by $DF_{\rho}$. Correlations are computed for all pairs of IDRs and PFAs across both structural archetypes.
@@ 3.5 Metric 5: Sensitivity indices to aleatory and epistemic uncertainties
Variance-based sensitivity indices, computed using Saltelli’s method [40] (details in B), quantify the contributions of aleatory (ground motion) and epistemic (structural parameter) uncertainties to seismic responses. This group-wise sensitivity analysis is conducted across the four cases defined in Table 2, examining how variations in structural parameter distributions affect sensitivity patterns.
@@
For the long-period tower, sensitivity indices and their discrepancies, shown in Figures 13 and 14, exhibit trends similar to those observed for the medium-period building, with overall differences remaining below 30%. These results confirm the applicability of synthetic ground motions for analyzing global sensitivity, while emphasizing the growing influence of structural uncertainties in taller structures, particularly when the variability in uncertain parameters is large.
@@
$S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$
@@
$S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$ $S_{\mathbf{S}}$ $S_{\mathbf{GM}}$
@@ 4 Synthesis of the comparative analysis
    Significant discrepancies in extreme behaviors: Metrics associated with extreme behaviors, such as tail indices and higher-order normalized moments (skewness and kurtosis), exhibit pronounced discrepancies, often exceeding 50%. These differences highlight the limitations of synthetic motions in capturing the complex, non-Gaussian features of recorded motions, including pulse-like effects, extreme variability, and other rare-event phenomena. These findings underscore the need for caution when relying solely on synthetic motions to assess extreme responses or rare-event probabilities.
@@
    Relative importance of aleatory and epistemic uncertainties: Sensitivity analysis shows that ground motion variability (aleatory uncertainty) consistently dominates structural parameter variability (epistemic uncertainty) in driving seismic responses. However, for both structural models, the relative influence of structural uncertainties increases with building height, particularly for IDRs in upper stories, where structural variability can surpass ground motion variability in some cases. This effect is most pronounced when structural variability is high but remains notable across all studied cases with varying distribution types. In contrast, for PFAs, ground motion variability remains the dominant factor across all cases, regardless of structural variability. These sensitivity patterns are consistent across both synthetic and recorded datasets, confirming the applicability of synthetic motions for capturing global sensitivity trends, while emphasizing the growing influence of structural uncertainties in taller structures with large parameter variability.
@@ 5 Conclusions
This study presents a comparative analysis of stochastic seismic responses under synthetic and recorded ground motions, emphasizing their implications for structural design and seismic risk assessment. Five key metrics—probability distributions, statistical moments, correlations, tail indices, and sensitivity indices—are evaluated using two archetypal structures: a 12-story medium-period building and a high-rise long-period tower. Both ground motion datasets are calibrated to a shared target spectrum, ensuring consistency in spectral characteristics and enabling a robust comparison of their effects on structural responses.
Synthetic motions demonstrate reliability in capturing global response characteristics. Differences in distributions, mean and variance, correlations, and sensitivity indices generally remain within 10%–20%, validating their suitability for routine seismic design and global probabilistic response modeling. These findings underscore the value of synthetic motions as efficient tools for representing central tendencies and overall variability in seismic responses. However, significant discrepancies arise in metrics related to extreme behaviors, particularly for long-period structures. Differences in tail indices, higher-order moments, and responses sensitive to low-frequency components frequently exceed 50%. These discrepancies highlight the limitations of synthetic motions in replicating the complex, non-Gaussian features of recorded motions, such as pulse-like effects and extreme variability—limitations that are critical for rare-event analysis and the design of tall buildings or other long-period structures. Sensitivity analysis further reveals that ground motion variability generally dominates structural parameter variability in driving seismic responses, with consistent trends observed across both synthetic and recorded ground motions.
In conclusion, when carefully calibrated, synthetic ground motions provide an efficient and reliable means of capturing global seismic responses, making them suitable for routine seismic analyses. However, recorded ground motions remain indispensable for applications requiring accurate representation of rare-event scenarios or dynamic phenomena critical to long-period structures. Future research could focus on enhancing stochastic ground motion models to better replicate non-Gaussian features and complex earthquake phenomena, bridging the gap between computational efficiency and real-world accuracy in seismic design and risk assessment.
@@ Appendix A Algorithms for ground motion selection and generation
This study employs recorded and synthetic ground motion datasets that are calibrated to match a shared target spectrum, ensuring consistency in spectral median, variance, and correlation structure. The datasets are constructed using two advanced algorithms: a ground motion selection method [17] for recorded motions and a ground motion generation method [19] for synthetic motions. The procedures are as follows, with both methods sharing the first two steps:
@@
      For recorded motions: The selection algorithm described in [17] is employed. This approach selects $N_{t}$ ground motions from a candidate dataset based on their similarity to the target spectrum. The process involves computing error metrics between the candidate motion spectra and each target spectrum realization, selecting those with the smallest error, and optionally scaling their amplitudes for better spectral matching.
@@
      For synthetic motions: The generation algorithm described in [19] is used. This method employs spectral representation techniques [41] with time-frequency modulating function to generate artificial ground motions. Iterative adjustments are made to align the generated motions with the target spectrum. A baseline correction ensures realistic time histories, resulting in non-stationary accelerograms compatible with the target spectrum.
@@ Appendix B Computation of sensitivity indices using Saltelli’s method
Variance-based sensitivity indices are critical for quantifying the contributions of input uncertainties to seismic responses. However, their computation can be computationally intensive, particularly for high-dimensional models frequently encountered in seismic response analyses. Traditional variance decomposition methods often require a double-loop integration process, leading to a substantial number of model evaluations. To address this challenge, Saltelli’s scheme [40] is utilized in this study. This method employs stratified quasi-random sampling to compute sensitivity indices efficiently, significantly reducing computational costs, and has been widely applied in engineering domains [42, 43, 44].

### v2

@@ 3.1 Metric 1: Probability distributions
Figures 6 and 7 present the empirical PDFs and complementary cumulative distribution functions (CCDFs) of EDPs—peak IDRs and PFAs—for the medium-period building and the long-period tower. Both structural models exhibit close agreement between responses to synthetic and recorded ground motions in the central ranges of the PDFs, indicating comparable global response characteristics under typical seismic scenarios. This alignment suggests that calibrating the synthetic and recorded ground motions to a target spectrum ensures consistency in global response characteristics. However, discrepancies become apparent in the tails of the distributions, particularly for the IDRs of the long-period tower. These differences can be attributed to the non-bijective mapping between spectral acceleration and ground motions, where both selection and generation methods introduce variability in time-domain characteristics despite comparable response spectra.
@@ 3.2 Metric 2: Statistical moments
Figure 9 illustrates differences in normalized moments ($DF_{M_{k}}$, i=k,,4i=k,...,4)—mean, variance, skewness, and kurtosis—between responses to synthetic and recorded ground motions for both structural models. Differences in mean and variance remain within 3%–20%, indicating close agreement in central tendencies and overall variability. These results demonstrate that well-calibrated synthetic motions can effectively capture general response characteristics.
@@ 3.3 Metric 3: Correlations between EDPs
Figure 10 presents correlation matrices for EDPs under synthetic and recorded motions, along with the corresponding differences quantified by $DF_{\rho}$. Correlations are computed for all pairs of IDRs and PFAs across both structural models.
@@ 3.5 Metric 5: Sensitivity indices to aleatory and epistemic uncertainties
Variance-based sensitivity indices, computed using Saltelli’s method saltelli2010variance (details in B), quantify the contributions of aleatory (ground motion) and epistemic (structural parameter) uncertainties to seismic responses. This group-wise sensitivity analysis is conducted across the four cases defined in Table 2, examining how variations in structural parameter distributions affect sensitivity patterns.
@@
For the long-period tower, sensitivity indices and their discrepancies, shown in Figures 14 and 15, exhibit trends similar to those observed for the medium-period building, with overall differences remaining below 30%. These results confirm the applicability of synthetic ground motions for analyzing global sensitivity, while emphasizing the growing influence of structural uncertainties in taller structures, particularly when the variability in uncertain parameters is large.
@@
$S_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $S_{\mathbf{GM}}$
@@
$S_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $_{\mathbf{GM}} S𝐒$$_{\mathbf{S}}$ $S_{\mathbf{GM}}$
@@ 4 Synthesis of the comparative analysis
    Significant discrepancies in extreme behaviors: Metrics associated with extreme behaviors, such as tail indices and higher-order normalized moments (skewness and kurtosis), exhibit pronounced discrepancies, often exceeding 50%. These differences are particularly prominent for story drift responses in long-period structures. Synthetic motions consistently underpredict the frequency and severity of large excursions, a limitation that cannot be identified by examining only low-order statistics. Most previous studies have validated synthetic motions primarily using spectral compatibility or mean/variance agreement, which overlook these tail discrepancies. This study reveals that such underestimation of extremes poses non-conservative risks, particularly in performance-based seismic design, where rare events govern decision-making. These results highlight that synthetic ground motions—while useful for average behavior—may not be suitable for simulating extreme responses without further refinement.
@@
    Relative importance of aleatory and epistemic uncertainties: Sensitivity analysis shows that ground motion variability consistently dominates structural parameter variability in driving seismic responses. However, for both structural models, the relative influence of structural uncertainties increases with building height, particularly for IDRs in upper stories, where structural variability can surpass ground motion variability in some cases. This effect is most pronounced when structural variability is high but remains notable across all studied cases with varying distribution types. In contrast, for PFAs, ground motion variability remains the dominant factor across all cases, regardless of structural variability. These sensitivity patterns are consistent across both synthetic and recorded datasets, confirming the applicability of synthetic motions for capturing global sensitivity trends, while emphasizing the growing influence of structural uncertainties in taller structures with large parameter variability.
@@ 5 Conclusions
This study presents a detailed comparative analysis of stochastic seismic responses under synthetic and recorded ground motions, focusing on five key metrics: probability distributions, statistical moments, correlations, tail indices, and sensitivity indices. Both ground motion datasets are calibrated to a shared target spectrum, ensuring consistency in spectral characteristics and enabling a robust comparison of their effects on structural responses. Two structural models—a 12-story medium-period building and a 110-story long-period tower—are used to investigate how these response metrics vary with ground motion type across different dynamic regimes.
The results show that synthetic motions are reliable surrogates for recorded motions in capturing global response trends. Across both structural models, differences in distributions, mean and variance, correlations, and variance-based sensitivity indices are generally small (within 10%–20%), validating their suitability for routine seismic design and global probabilistic response modeling. These findings underscore the value of synthetic motions as efficient tools for representing central tendencies and overall variability in seismic responses. However, significant discrepancies arise in metrics related to extreme behaviors, particularly for long-period structures. Metrics associated with extreme events, including the tails of story drift distributions and higher-order moments, exhibit large discrepancies, often exceeding 50%. These discrepancies highlight the limitations of synthetic motions in replicating the complex, non-Gaussian features of recorded motions, such as pulse-like effects and extreme variability—limitations that are critical for rare-event analysis and the design of tall buildings or other long-period structures. This underscores the need for caution when using synthetic motions in rare-event risk assessments or design of tall buildings.
In conclusion, when carefully calibrated, synthetic motions provide a computationally efficient and statistically consistent means of capturing global seismic response characteristics, but their use remains limited for applications requiring accurate representation of rare events or long-period structural responses. This distinction is vital: relying solely on synthetic motions without addressing tail behavior may compromise the reliability of rare-event analyses and performance-based seismic design.
[…]
@@ Acknowledgments
Data availability
The source codes are available for download at https://github.com/Jungh0Kim/GM_Comparative.
@@ Appendix A Algorithms for ground motion selection and generation
This study employs recorded and synthetic ground motion datasets that are calibrated to match a shared target spectrum, ensuring consistency in spectral median, variance, and correlation structure. The datasets are constructed using two advanced algorithms: a ground motion selection method baker2018improved for recorded motions and a ground motion generation method yanni2024probabilistic for synthetic motions. The procedures are as follows, with both methods sharing the first two steps:
@@
      For recorded motions: The selection algorithm described in baker2018improved is employed. This approach selects $N_{t}$ ground motions from a candidate dataset based on their similarity to the target spectrum. The process involves computing error metrics between the candidate motion spectra and each target spectrum realization, selecting those with the smallest error, and optionally scaling their amplitudes for better spectral matching.
@@
      For synthetic motions: The generation algorithm described in yanni2024probabilistic is used. This method employs spectral representation techniques vanmarcke1977simulated with time-frequency modulating function to generate artificial ground motions. Iterative adjustments are made to align the generated motions with the target spectrum. A baseline correction ensures realistic time histories, resulting in non-stationary accelerograms compatible with the target spectrum.
@@ Appendix B Computation of sensitivity indices using Saltelli’s method
Variance-based sensitivity indices are critical for quantifying the contributions of input uncertainties to seismic responses. However, their computation can be computationally intensive, particularly for high-dimensional models frequently encountered in seismic response analyses. Traditional variance decomposition methods often require a double-loop integration process, leading to a substantial number of model evaluations. To address this challenge, Saltelli’s scheme saltelli2010variance is utilized in this study. This method employs stratified quasi-random sampling to compute sensitivity indices efficiently, significantly reducing computational costs, and has been widely applied in engineering domains shin2013addressing , mahmoudi2019global , dela2022multi .